# On Edge Coloring of Multigraphs

### v4

@@ 3 Extented Tashkinov Tree (ETT)
    If $\eta\notin\overline{\varphi}_{1}(T^{1})$ and $P_{1}$ has an $(\eta,\tau)$-exit vertex of $T^{1}$ in $P_{1}(rw)-\{w\}$, then $\alpha\notin\overline{\varphi}_{1}(T^{1})$ and $T^{1}$ has a $(\pi,\alpha)$-exit vertex $x_{2}$ in $P_{z}(\pi,\alpha,\varphi_{1})$ with $x\prec x_{2}$ along $T^{1}$, and the corresponding $(\pi,\alpha)$-exit path does not share vertex with $T_{n+1}$.
Clearly $\varphi_{1}^{\prime}=\varphi/P_{x}(\alpha,\delta,\varphi)$ is $(T_{j}(x)-x,D_{j-1},\varphi)$-stable, because $P_{x}(\alpha,\delta,\varphi)\cap T_{n+1}=\{x\}$. Moreover, since $T_{n+1}$ is $RE$ finished, all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi$ are contained in $G[V(T_{n+1})]$ following Algorithm 3.1. So all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi$ are colored the same under $\varphi_{1}^{\prime}$ if $h\geq 1$. As $T_{n+1}$ is closed for both $\alpha$ and $\pi$ under $\varphi$, all the $(\pi,\alpha)$-chains intersect $T_{n+1}$ under $\varphi_{1}^{\prime}$ are contained in $G[V(T_{n+1})]$ except for $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ which contains $x$ who is the only $(\pi,\alpha)$-exit of $T_{n+1}$ under $\varphi_{1}^{\prime}$, and all vertices in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ from $z$ to $x$ are contained in $T_{n+1}$. So none of the $(\pi,\alpha)$-chains we switched by Kempe changes from $\varphi_{1}^{\prime}$ to $\varphi_{1}$ intersect $T_{n+1}$ and none of the $\alpha$ colored edges of $P\cup P_{ex}\cup D$ under $\varphi_{1}^{\prime}$ is in the subpath of $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ from $z$ to $x$. Thus we have $(a_{1})$. Moreover, since $\pi\notin\varphi(P\cup P_{ex}\cup D)$ by the assumptions in the statement of the current Lemma, we have $(e_{1})$.
By $(a_{1})$ and (2), we see that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{1}$ as under $\varphi$ if $h\geq 1$. Since $\varphi_{1}$ is $(T_{j}(x)-x,D_{j-1},\varphi)$-stable and $x\notin T_{j-1}$, we see that $\varphi_{1}$ is also $(T_{j-1},D_{j-1},\varphi)$-stable, and thus is $T_{j-1}+f_{j-1}$ mod $\varphi$ following (A2.5), and therefore $T_{j}(x)$ is an ETT with $j-1$-rungs under $\varphi_{1}$ following Corollary 3.1. Since $T^{1}$ is a closure of $T_{j}(x)$ under $\varphi_{1}$ and $j-1\leq n$, we have $(b_{1})$ by (A1) on $T^{1}$.
@@
Note that if $\eta\notin\overline{\varphi}_{1}(T^{1})$, then because $T^{1}$ is elementary by $(b_{1})$, we see that $P_{1}$ must have an $(\eta,\tau)$-exit vertex of $T^{1}$ following $(c_{1})$. To prove $(f_{1})$, we may assume that $\eta\notin\overline{\varphi}_{1}(T^{1})$ and $P_{1}$ has an $(\eta,\tau)$-exit vertex of $T^{1}$ in $P_{1}(rw)-\{w\}$. By $(iv)$ of $P$, (4), and $(c_{1})$, all the $\eta$ colored edges of $P_{2}$ and $P_{3}$ (see (3)) are in $P_{1}$ after $w$ along the direction from $r$. Because the $(\eta,\tau)$-exit vertex of $T^{1}$ is in $P_{1}(rw)-\{w\}$, we see that
@@
Following $(c_{1})$, we have $\varphi_{1}(P_{2}\cup P_{3})\subseteq(\overline{\varphi}_{1}(T^{1})\cup\{\alpha,\eta\}$. We first assume otherwise that $\alpha\in\overline{\varphi}_{1}(T^{1})$. Then by (5) and TAA, we see that no edges of $P_{2}$ and $P_{3}$ are colored by $\eta$ and $u,v\in V(T^{1})$. Since $T^{1}$ is elementary by $(b_{1})$, we have $\eta\in(\overline{\varphi}_{1}(u)\cup\overline{\varphi}_{1}(v))\subseteq\overline{\varphi}_{1}(T^{1})$ by $(d_{1})$, a contradiction. So $\alpha\notin\overline{\varphi}_{1}(T^{1})$, and therefore $T^{1}$ has a $(\alpha,\pi)$-exit vertex in $P_{z}(\pi,\alpha,\varphi_{1})$, as $T^{1}$ is elementary by $(b_{1})$. Moreover, none of $u$ and $v$ is in $T^{1}$, as otherwise we similarly have $\alpha\in\overline{\varphi}_{1}(T^{1})$ by $(c_{1})$ and the assumption that $\eta\notin\overline{\varphi}_{1}(T^{1})$. So following (5) and TAA, the two edge disjoint paths $P_{2}$ and $P_{3}$ contain two $\alpha$ colored boundary edges $f$ and $g$ of $T^{1}$. Since all the $\alpha$ colored edges of $P_{2}$ and $P_{3}$ belong to $P_{z}(\pi,\alpha,\varphi_{1})$ after the vertex $x$ along the direction from $z$ by $(e_{1})$, one of $f$ and $g$ is incident to a vertex in $T^{1}$ other than $x$ and it is after $x$ along $P_{z}(\pi,\alpha,\varphi_{1})$ along the direction from $z$. Since $T^{1}$ has a $(\alpha,\pi)$-exit vertex in $P_{z}(\pi,\alpha,\varphi_{1})$, we see that $T^{1}$ has a $(\pi,\alpha)$-exit vertex $x_{2}$ in $P_{z}(\pi,\alpha,\varphi_{1})$ where $x_{2}$ is after $x$ along $P_{z}(\pi,\alpha,\varphi_{1})$ from the direction starting at $z$. Because $P_{z}(\pi,\alpha,\varphi_{1})$ and $T_{n+1}$ only share vertices on the subpath of $P_{z}(\pi,\alpha,\varphi_{1})$ from $z$ to $x$ (which is $P_{z}(\pi,\alpha,\varphi)$) by the construction of $\varphi_{1}$ ($T_{n+1}$ is closed for $\pi$ and it only has one boundary edge colored by $\alpha$ which is incident to $x$ under $\varphi_{1}$), we see that $x_{2}\notin T_{n+1}$ and the corresponding $(\pi,\alpha)$-exit path of $T^{1}$ under $\varphi_{1}$ does not share vertex with $T_{n+1}$. So $x\prec x_{2}$ along $T^{1}$ and $(f_{1})$ is established.
@@
Now we have as claimed. Suppose $\varphi_{i-1}$ ($i\geq 2$) is constructed with claims $(a_{i-1})$ to $(f_{i-1})$ in place of index $i$ bellow (note that $\varphi_{1}$ also satisfies the claim below with $\alpha_{1}=\alpha$, $x_{1}=x$, $T^{0}=T_{j}$, and $\varphi_{0}=\varphi$). If $\eta\in\overline{\varphi}_{i-1}(T^{i-1})$ or $P_{1}$ does not contain an $(\eta,\tau)$-exit vertex of $T^{i-1}$ in $P_{1}(rw)-\{w\}$, we stop and let $\mu=\varphi_{i-1}$. Otherwise, we recursively construct $\varphi_{i}$ from $\varphi_{i-1}$ and $x_{i}$ with an arbitrary color $\alpha_{i}\in\overline{\varphi}_{i-1}(x_{i})$, and have the following claims. Here is the detailed construction: We first let $\varphi^{\prime}=\varphi_{i-1}/(G-T^{i-1},\alpha_{i},\pi)$, and then let $\varphi_{i}^{\prime}=\varphi^{\prime}/P_{x_{i}}(\alpha_{i},\alpha_{i-1},\varphi^{\prime})$. Finally, we construct $\varphi_{i}$ from $\varphi_{i}^{\prime}$ by switching all the $\alpha_{i}$ colored edges of $P\cup P_{ex}\cup D$ not in $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ to $\pi$ by Kempe changes. Let $T^{i}$ be a closure of $T^{i-1}(x_{i})$ under $\varphi_{i}$.
@@
    $T^{i}$ is an elementary ETT with the same $j-1$-rungs under $\varphi_{i}$ as $T_{j}$ under $\varphi$.
@@
    All the $\alpha_{i}$-colored edges of $P\cup P_{ex}\cup D$ under $\varphi_{i}$ belong to $P_{z}(\pi,\alpha_{i},\varphi_{i})$ after the vertex $x_{i}$ along the direction from $z$.
@@
    If $\eta\notin\overline{\varphi}_{i}(T^{i})$ and $P_{1}$ has an $(\eta,\tau)$-exit vertex of $T^{i}$ in $P_{1}(rw)-\{w\}$, then $\alpha_{i}\notin\overline{\varphi}_{i}(T^{i})$ and $T^{i}$ has a $(\pi,\alpha_{i})$-exit vertex $x_{i+1}$ in $P_{z}(\pi,\alpha_{i},\varphi_{i})$ with $x_{i}\prec x_{i+1}$ along $T^{i}$, and the corresponding $(\pi,\alpha_{i})$-exit path does not share vertex with $T^{i-1}$.
Now we prove the above claims inductively. By $(f_{k})$ for all $11\leq k\leq i-$ 1 and (1), we see that $x\prec x_{1},...,\prec x_{i}$ along $T^{i-1}$ and $x_{i}\notin T_{h}$. Following (4), $(a_{i-1})$, $(b_{i-1})$, $(c_{i-1})$, and the assumption of $\eta\notin\overline{\varphi}_{i-1}(T^{i-1})$ ($i\geq 2$) and the $(\eta,\tau)$-exit vertex of $T^{i-1}$ being in $P_{1}(rw)-\{w\}$, we have
(6) $\alpha_{i}\notin\{\eta,\alpha_{i-1}\}$ and thus $\pi,\alpha_{i},\eta,\alpha_{i-1}$, $\gamma_{h}$ (when $h\geq 1$) and $\tau$ are all different, and $\pi,\tau\notin S_{h}$ and $\alpha_{i}\notin\varphi_{i}(P\cup P_{ex}\cup D)$.
@@
By (A1), (A2), and $(b_{i-1})$ on $T_{h}$ under $\varphi_{i-1}$ , we have
@@
By $(a_{i})$ and (8), we see that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{i}$ as under $\varphi_{i-1}$ if $h\geq 1$. By $(f_{k})$ for all $11\leq k\leq i-$ 1, we have $x_{i}\notin T_{j-1}$. Since $\varphi_{i}$ is $(T^{i-1}(x_{i})-x_{i},D_{j-1},\varphi_{i-1})$-stable and $x_{i}\notin T_{j-1}$, we see that $\varphi_{i}$ is also $(T_{j-1},D_{j-1},\varphi_{i-1})$-stable ($T^{i-1}$ has the same rung under $\varphi_{i-1}$ as $T_{j}$ under $\varphi$ by $(b_{i-1})$), and thus is a $T_{j-1}+f_{j-1}$ mod $\varphi_{i-1}$ coloring by (A2.5). Therefore, following Corollary 3.1, $T^{i-1}(x_{i})$ is an ETT with the same $(j-1)$-rungs under $\varphi_{i}$ as $T^{i-1}$ under $\varphi_{i-1}$. Since $T^{i}$ is a closure of $T^{i-1}(x_{i})$ under $\varphi_{i}$ and $j-1\leq n$, we have $(b_{i})$ by (A1) on $T^{i}$ and $(b_{i-1})$.
The rest parts of the claim above are essentially the same as in the proof of the earlier claim with changed index, here we omit the details.
Now with the above claim, we see that we can always recursively construct $\varphi_{i}$ with a new vertex $x_{i+1}$ if $\eta\notin\overline{\varphi}_{i-1}(T^{i-1})$ ($i\geq 2$) and the $(\eta,\tau)$-exit vertex of $T^{i-1}$ is in $P_{1}(rw)-\{w\}$. But since we have a finite graph, at a certain point we have to stop at some $\varphi_{i}$ when $\eta\in\overline{\varphi}_{i}(T^{i})$ or the $(\eta,\tau)$-exit vertex of $T^{i}$ is in $P_{1}(wr^{\prime})$; we still denote the corresponding coloring $\varphi_{i}$ by $\mu$. So $\mu$ satisfies all conditions $(a_{i})$ to $(f_{i})$.
Note that by $(a_{i})$, $(c_{i})$, the fact that $x\prec x_{i}$ (see the paragraph above (6)) and the assumption $(iv)$ of $P$, we have
(9) $\tau\in\overline{\mu}(z)$ and $\tau\in\overline{\mu}(T^{i})$.
We first consider the case that $\eta\in\overline{\mu}(T^{i})$. Then following assumption $(iv)$ of $P$, we see that $w\in V(T^{i})$ by TAA. So $\alpha_{i}\in\overline{\mu}(w)$ following $(d_{i})$, because $T^{i}$ is elementary by $(b_{i})$. By $(c_{i})$, both $u$ and $v$ are in $T_{n+1}$ by TAA. However, following $(d_{i})$, we reach a contradiction to the elementariness of $T_{n+1}$. So we may assume that
(10) $\eta\notin\overline{\mu}(T^{i})$, and $P_{1}$ has an $(\eta,\tau)$-exit vertex $y_{1}$ of $T^{i}$ in $P_{1}(wr^{\prime})$.
@@
Let $\mu_{0}=\mu$, $y_{0}=x_{i}$, $\beta_{0}=\eta$ and $T^{0}_{\mu}=T^{i}$. Suppose we have already constructed $\mu_{l-1}$ and $T^{l-1}_{\mu}$, and suppose they satisfy the claims bellow (with index $l-1$) if $l\geq 2$. If $\alpha_{i}\in\overline{\mu}_{l-1}(T^{l-1}_{\mu}(y_{l}))$, we stop. Otherwise, we pick an arbitrary color $\beta_{l}$ in $\overline{\mu}_{l-1}(y_{l})$ and let $\mu^{\prime}=\mu_{l-1}/(G-T^{l-1}_{\mu},\beta_{l},\tau)$, and then let $\mu_{l}^{\prime}=\mu^{\prime}/P_{y_{l}}(\beta_{l},\beta_{l-1},\mu^{\prime})$. Finally, we construct $\mu_{l}$ from $\mu_{l}^{\prime}$ by switching all the $\beta_{l}$ colored edges of $P\cup P_{ex}\cup D\cup P_{1}$ not in $P_{r}(\tau,\beta_{l},\mu_{l}^{\prime})$ to $\tau$ by Kempe changes. Let $T^{l}_{\mu}$ be a closure of $T^{l-1}_{\mu}(y_{l})$ under $\mu_{i}$.
@@
    $T^{l}_{\mu}$ is an elementary ETT with the same $j-1$-rungs under $\mu_{l}$ as $T^{i}$ under $\mu$.
@@
    All the $\beta_{l}$-colored edges of $P\cup P_{ex}\cup D\cup P_{1}$ under $\mu_{l}$ belong to $P_{r}(\tau,\beta_{l},\mu_{l})$ after the vertex $y_{l}$ along the direction from $r$.
@@
We then prove the above claims inductively in a similar fashion as earlier. Here are the details. Recall that $x\prec x_{i}$ (see the paragraph above (6)). By $(F_{k})$ for all $11\leq k\leq l-$ 1, (11) and (1), we see that
(12) $x_{i}=y_{0}\prec y_{1},...,\prec y_{l}$ along $T^{i-1}_{\mu}$ and $y_{l}\notin T_{h}$.

### v5

@@ 3 Extented Tashkinov Tree (ETT)
    If $\eta\notin\overline{\varphi}_{1}(T^{1})$ and $P_{1}$ has an $(\eta,\tau)$-exit vertex of $T^{1}$ in $P_{1}[r,w]-\{w\}$, then $\alpha\notin\overline{\varphi}_{1}(T^{1})$ and $T^{1}$ has a $(\pi,\alpha)$-exit vertex $x_{2}$ in $P_{z}(\pi,\alpha,\varphi_{1})$ with $x\prec x_{2}$ along $T^{1}$, and the corresponding $(\pi,\alpha)$-exit path does not share vertex with $T_{n+1}$.
Clearly $\varphi_{1}^{\prime}=\varphi/P_{x}(\alpha,\delta,\varphi)$ is $(T_{j}(x)-x,D_{j-1},\varphi)$-stable, because $P_{x}(\alpha,\delta,\varphi)\cap T_{n+1}=\{x\}$. Moreover, since $T_{n+1}$ is $\operatorname{RE}$ finished, all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi$ are contained in $G[V(T_{n+1})]$ following Algorithm 3.1. So all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi$ are colored the same under $\varphi_{1}^{\prime}$ if $h\geq 1$. As $T_{n+1}$ is closed for both $\alpha$ and $\pi$ under $\varphi$, all the $(\pi,\alpha)$-chains intersect $T_{n+1}$ under $\varphi_{1}^{\prime}$ are contained in $G[V(T_{n+1})]$ except for $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ which contains $x$ who is the only $(\pi,\alpha)$-exit of $T_{n+1}$ under $\varphi_{1}^{\prime}$, and all vertices in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ from $z$ to $x$ are contained in $T_{n+1}$. So none of the $(\pi,\alpha)$-chains we switched by Kempe changes from $\varphi_{1}^{\prime}$ to $\varphi_{1}$ intersect $T_{n+1}$ and none of the $\alpha$ colored edges of $P\cup P_{ex}\cup D$ under $\varphi_{1}^{\prime}$ is in the subpath of $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ from $z$ to $x$. Thus we have $(a_{1})$. Moreover, since $\pi\notin\varphi(P\cup P_{ex}\cup D)$ by the assumptions in the statement of the current Lemma, we have $(e_{1})$.
By $(a_{1})$ and (2), we see that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{1}$ as under $\varphi$ if $h\geq 1$. Since $\varphi_{1}$ is $(T_{j}(x)-x,D_{j-1},\varphi)$-stable and $x\notin T_{j-1}$, we see that $\varphi_{1}$ is also $(T_{j-1},D_{j-1},\varphi)$-stable, and thus is $T_{j-1}+f_{j-1}$ mod $\varphi$ following (A2.5), and therefore $T_{j}(x)$ is an ETT with j1j-1-rungs under $\varphi_{1}$ following Corollary 3.1. Since $T^{1}$ is a closure of $T_{j}(x)$ under $\varphi_{1}$ and $j-1\leq n$, we have $(b_{1})$ by (A1) on $T^{1}$.
@@
Note that if $\eta\notin\overline{\varphi}_{1}(T^{1})$, then because $T^{1}$ is elementary by $(b_{1})$, we see that $P_{1}$ must have an $(\eta,\tau)$-exit vertex of $T^{1}$ following $(c_{1})$. To prove $(f_{1})$, we may assume that $\eta\notin\overline{\varphi}_{1}(T^{1})$ and $P_{1}$ has an $(\eta,\tau)$-exit vertex of $T^{1}$ in $P_{1}[r,w]-\{w\}$. Since $\varphi(P_{ex}\cup D)\subseteq\{\alpha,\delta\}$, by $(iv)$ of $P$, (4), and $(c_{1})$, all the $\eta$ colored edges of $P_{2}$ and $P_{3}$ (see (3)) are in $P_{1}$ after $w$ along the direction from $r$. Because the $(\eta,\tau)$-exit vertex of $T^{1}$ is in $P_{1}[r,w]-\{w\}$, we see that
@@
Following $(c_{1})$, we have $\varphi_{1}(P_{2}\cup P_{3})\subseteq\overline{\varphi}_{1}(T^{1})\cup\{\alpha,\eta\}$. We first assume otherwise that $\alpha\in\overline{\varphi}_{1}(T^{1})$. Then by (5) and TAA, we see that no edges of $P_{2}$ and $P_{3}$ are colored by $\eta$ and $u,v\in V(T^{1})$. Since $T^{1}$ is elementary and $T^{1}(x)=T_{j}(x)\subseteq T_{n+1}$ by $(b_{1})$, we have $\eta\in(\overline{\varphi}_{1}(u)\cup\overline{\varphi}_{1}(v))\subseteq\overline{\varphi}_{1}(T^{1})$ by $(d_{1})$ and the fact that $u,v\notin V(T_{n+1})$, a contradiction to the assumption that $\eta\notin\overline{\varphi}_{1}(T^{1})$. So $\alpha\notin\overline{\varphi}_{1}(T^{1})$, and therefore $T^{1}$ has a $(\alpha,\pi)$-exit vertex in $P_{z}(\pi,\alpha,\varphi_{1})$, as $T^{1}$ is elementary by $(b_{1})$. Moreover, none of $u$ and $v$ is in $T^{1}$, as otherwise we similarly have $\alpha\in\overline{\varphi}_{1}(T^{1})$ by $(c_{1})$ and the assumption that $\eta\notin\overline{\varphi}_{1}(T^{1})$. So following (5) and TAA, the two edge disjoint paths $P_{2}$ and $P_{3}$ contain two $\alpha$ colored boundary edges $f$ and $g$ of $T^{1}$. Since all the $\alpha$ colored edges of $P_{2}$ and $P_{3}$ belong to $P_{z}(\pi,\alpha,\varphi_{1})$ after the vertex $x$ along the direction from $z$ by $(e_{1})$, one of $f$ and $g$ is incident to a vertex in $T^{1}$ other than $x$ and it is after $x$ along $P_{z}(\pi,\alpha,\varphi_{1})$ along the direction from $z$. Since $T^{1}$ has an $(\alpha,\pi)$-exit vertex in $P_{z}(\pi,\alpha,\varphi_{1})$, we see that $T^{1}$ has a $(\pi,\alpha)$-exit vertex $x_{2}$ in $P_{z}(\pi,\alpha,\varphi_{1})$ where $x_{2}$ is after $x$ along $P_{z}(\pi,\alpha,\varphi_{1})$ from the direction starting at $z$. Because $P_{z}(\pi,\alpha,\varphi_{1})$ and $T_{n+1}$ only share vertices on the subpath of $P_{z}(\pi,\alpha,\varphi_{1})$ from $z$ to $x$ (which is $P_{z}(\pi,\alpha,\varphi)$) by the construction of $\varphi_{1}$ ($T_{n+1}$ is closed for $\pi$ and it only has one boundary edge colored by $\alpha$ which is incident to $x$ under $\varphi_{1}$), we see that $x_{2}\notin T_{n+1}$ and the corresponding $(\pi,\alpha)$-exit path of $T^{1}$ under $\varphi_{1}$ does not share vertex with $T_{n+1}$. So $x\prec x_{2}$ along $T^{1}$ and $(f_{1})$ is established.
@@
Now we have as claimed. Suppose $\varphi_{i-1}$ ($i\geq 2$) is constructed with claims $(a_{i-1})$ to $(f_{i-1})$ bellow (note that $\varphi_{1}$ also satisfies the claims below with $\alpha_{1}=\alpha$, $x_{1}=x$, $T^{0}=T_{j}$, and $\varphi_{0}=\varphi$). If $\eta\in\overline{\varphi}_{i-1}(T^{i-1})$ or $P_{1}$ does not contain an $(\eta,\tau)$-exit vertex of $T^{i-1}$ in $P_{1}[r,w]-\{w\}$, we stop and let $\mu=\varphi_{i-1}$. Otherwise, we recursively construct $\varphi_{i}$ from $\varphi_{i-1}$ and $x_{i}$ with an arbitrary color $\alpha_{i}\in\overline{\varphi}_{i-1}(x_{i})$, and have the following claims. Here is the detailed construction: We first let $\varphi^{\prime}=\varphi_{i-1}/(G-T^{i-1},\alpha_{i},\pi)$, and then let $\varphi_{i}^{\prime}=\varphi^{\prime}/P_{x_{i}}(\alpha_{i},\alpha_{i-1},\varphi^{\prime})$. Finally, we construct $\varphi_{i}$ from $\varphi_{i}^{\prime}$ by switching all the $\alpha_{i}$ colored edges of $P\cup P_{ex}\cup D$ not in $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ to $\pi$ by Kempe changes. Let $T^{i}$ be a closure of $T^{i-1}(x_{i})$ under $\varphi_{i}$.
@@
    $T^{i}$ is an elementary ETT with the same j1j-1-rungs under $\varphi_{i}$ as $T_{j}$ under $\varphi$.
@@
    All the $\alpha_{i}$ colored edges of $P\cup P_{ex}\cup D$ under $\varphi_{i}$ belong to $P_{z}(\pi,\alpha_{i},\varphi_{i})$ after the vertex $x_{i}$ along the direction from $z$.
@@
    If $\eta\notin\overline{\varphi}_{i}(T^{i})$ and $P_{1}$ has an $(\eta,\tau)$-exit vertex of $T^{i}$ in $P_{1}[r,w]-\{w\}$, then $\alpha_{i}\notin\overline{\varphi}_{i}(T^{i})$ and $T^{i}$ has a $(\pi,\alpha_{i})$-exit vertex $x_{i+1}$ in $P_{z}(\pi,\alpha_{i},\varphi_{i})$ with $x_{i}\prec x_{i+1}$ along $T^{i}$, and the corresponding $(\pi,\alpha_{i})$-exit path does not share vertex with $T^{i-1}$.
Now we prove the above claims inductively. By $(f_{k})$ for all $1\leq k\leq i-1$ and (1), we see that $x\prec x_{1}\prec...\prec x_{i}$ along $T^{i-1}$, and therefore $x_{i}\notin T_{h}$ by (1) and $(b_{i-1})$. Following (4), $(a_{i-1})$, $(b_{i-1})$, $(c_{i-1})$, and the assumption of $\eta\notin\overline{\varphi}_{i-1}(T^{i-1})$ ($i\geq 2$) and the $(\eta,\tau)$-exit vertex of $T^{i-1}$ being in $P_{1}[r,w]-\{w\}$, we have
(6) $\alpha_{i}\notin\{\eta,\alpha_{i-1}\}$, and thus $\pi,\alpha_{i},\eta,\alpha_{i-1}$, $\gamma_{h}$ (when $h\geq 1$), and $\tau$ are all different, and $\pi,\tau\notin S_{h}$.
@@
By (A2) and $(b_{i-1})$ under $\varphi_{i-1}$ , we have
@@
By $(a_{i})$ and (8), we see that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{i}$ as under $\varphi_{i-1}$ if $h\geq 1$. Note that $T^{i-1}$ has the same rung under $\varphi_{i-1}$ as $T_{j}$ under $\varphi$ following $(b_{i-1})$. So by $(f_{k})$ for all $1\leq k\leq i-1$, we have $x_{i}\notin T_{j-1}$. Since $\varphi_{i}$ is $(T^{i-1}(x_{i})-x_{i},D_{j-1},\varphi_{i-1})$-stable and $x_{i}\notin T_{j-1}$, we see that $\varphi_{i}$ is also $(T_{j-1},D_{j-1},\varphi_{i-1})$-stable, and thus is a $T_{j-1}+f_{j-1}$ mod $\varphi_{i-1}$ coloring by (A2.5). Therefore, following Corollary 3.1, $T^{i-1}(x_{i})$ is an ETT with the same (j1)(j-1)-rungs under $\varphi_{i}$ as $T^{i-1}$ under $\varphi_{i-1}$. Since $T^{i}$ is a closure of $T^{i-1}(x_{i})$ under $\varphi_{i}$ and $j-1\leq n$, we have $(b_{i})$ by (A1) on $T^{i}$ and $(b_{i-1})$.
The rest part of the claims above are essentially the same as in the proof of the earlier claims of $i=1$ with changed index, here we omit the details.
Now with the above claims, we see that we can always recursively construct $\varphi_{i}$ with a new vertex $x_{i+1}$ if $\eta\notin\overline{\varphi}_{i-1}(T^{i-1})$ ($i\geq 2$) and the $(\eta,\tau)$-exit vertex of $T^{i-1}$ is in $P_{1}[r,w]-\{w\}$. But since we have a finite graph, at a certain point we have to stop at some $\varphi_{i}$ when $\eta\in\overline{\varphi}_{i}(T^{i})$ or the $(\eta,\tau)$-exit vertex of $T^{i}$ is in $P_{1}[w,r^{\prime}]$; we still denote the corresponding coloring $\varphi_{i}$ by $\mu$. So $\mu$ satisfies all conditions $(a_{i})$ to $(f_{i})$.
Note that by $(a_{i})$, the fact that $x\prec x_{i}$ (see the paragraph above (6)) and the assumption $(iv)$ of $P$, we have
(9) $\tau\in\overline{\mu}(r)$, and therefore $\tau\in\overline{\mu}(T^{i})$. Moreover, $\mu(P_{1})\subseteq\{\tau,\eta\}$.
We first consider the case that $\eta\in\overline{\mu}(T^{i})$. Then in view of (9), we see that $w\in V(T^{i})$ by TAA along $P_{1}$. So $\alpha_{i}\in\overline{\mu}(w)$ following $(d_{i})$, because $T^{i}$ is elementary by $(b_{i})$ and $w\notin V(T_{j}(x))=V(T^{i}(x))$. By $(c_{i})$, both $u$ and $v$ are in $T_{n+1}$ by TAA. However, following $(d_{i})$ and the fact $u,v\notin V(T_{j}(x))=V(T^{i}(x))$, we reach a contradiction to the elementariness of $T_{n+1}$. So we may assume that
(10) $\eta\notin\overline{\mu}(T^{i})$, and $P_{1}$ has an $(\eta,\tau)$-exit vertex $y_{1}$ of $T^{i}$ in $P_{1}[w,r^{\prime}]$.
@@
Let $\mu_{0}=\mu$, $y_{0}=x_{i}$, $\beta_{0}=\eta$ and $T^{0}_{\mu}=T^{i}$. Suppose we have already constructed $\mu_{l-1}$ and $T^{l-1}_{\mu}$, and suppose they satisfy the claims bellow (with index l1l-1) if $l\geq 2$. If $\alpha_{i}\in\overline{\mu}_{l-1}(T^{l-1}_{\mu}(y_{l}))$, we stop. Otherwise, we pick an arbitrary color $\beta_{l}$ in $\overline{\mu}_{l-1}(y_{l})$ and let $\mu^{\prime}=\mu_{l-1}/(G-T^{l-1}_{\mu},\beta_{l},\tau)$. Then let $\mu_{l}^{\prime}=\mu^{\prime}/P_{y_{l}}(\beta_{l},\beta_{l-1},\mu^{\prime})$. Finally, we construct $\mu_{l}$ from $\mu_{l}^{\prime}$ by switching all the $\beta_{l}$ colored edges of $P\cup P_{ex}\cup D\cup P_{1}$ not in $P_{r}(\tau,\beta_{l},\mu_{l}^{\prime})$ to $\tau$ by Kempe changes. Let $T^{l}_{\mu}$ be a closure of $T^{l-1}_{\mu}(y_{l})$ under $\mu_{i}$.
@@
    $T^{l}_{\mu}$ is an elementary ETT with the same j1j-1-rungs under $\mu_{l}$ as $T^{i}$ under $\mu$.
@@
    All the $\beta_{l}$ colored edges of $P\cup P_{ex}\cup D\cup P_{1}$ under $\mu_{l}$ belong to $P_{r}(\tau,\beta_{l},\mu_{l})$ after the vertex $y_{l}$ along the direction from $r$.
@@
We then prove the above claims inductively in a similar fashion as earlier. Here are the details. Recall that $x\prec x_{i}$ (see the paragraph above (6)). By $(F_{k})$ for all $1\leq k\leq l-1$, (11), $(b_{i})$, $(B_{l-1})$ and (1), we see that
(12) $x_{i}=y_{0}\prec y_{1}\prec...\prec y_{l}$ along $T^{i-1}_{\mu}$ and $y_{l}\notin T_{h}$.